### v1

# High-charge relativistic electrons by vacuum laser acceleration from plasma mirrors using flying focus pulses

Abstract
Relativistic electron beams produced by intense lasers over short distances have important applications in high energy density physics and medical technologies. Vacuum laser acceleration with plasma mirrors injectors has garnered substantial research interest recently. However, a persistent challenge remains unresolved that electrons inevitably detach from the laser acceleration phase due to velocity mismatch. Here, we employ flying focus lasers to address this limitation. Through three-dimensional particle-in-cell simulations, we demonstrate that flying focus lasers can achieve a substantial enhancement in relativistic electron charge yield compared to conventional Gaussian lasers. This improvement stems from two key attributes: (1) The subluminal propagation velocity of the peak intensity keeps a larger electron population synchronized within the longitudinal ponderomotive acceleration region, and (2) Flying focus lasers sustain higher magnitudes of the longitudinal ponderomotive force over longer distances in comparison to Gaussian lasers. This approach offers high-charge relativistic electron sources ideal for demanding applications such as high-flux Thomson scattering and radiography.
I Introduction
The development of ultraintense femtosecond lasers Yu et al. (2012); Yoon et al. (2021) has enabled electron acceleration to relativistic energies over very short distance Esarey et al. (2009), paving the way for compact accelerator technologies with broad applications including ultrafast electron diffraction Morimoto and Baum (2018); He et al. (2016), fast ignition Tabak et al. (2005) and radiation therapy Fuchs et al. (2009). Over the past decades, two primary schemes have been extensively investigated both theoretically and experimentally: (1) Laser wakefield acceleration, in which intense lasers drive plasma wakefields in underdense plasmas, creating accelerating gradients typically on the order of 100 GV/m that can energize electrons to several GeV Tajima and Dawson (1979); Gonsalves et al. (2019); and (2) Vacuum laser acceleration (VLA), where electrons are directly accelerated by the intense fields provided by the laser pulses with gradients exceeding 10 TV/m Esarey et al. (1995); Malka et al. (1997); Varin and Piché (2006); De Andres et al. (2024).
In VLA, electron acceleration occurs in a vacuum environment where plasma field effects are negligible. This simplified mechanism has attracted significant efforts to enhance the accelerated beam qualities. However, electron injection in VLA is a critical challenge. For effective energy transfer from the laser field to electrons, two key prerequisites must be met: (1) an optimal injection position to align with the accelerating phase of the laser field, and (2) an initial injection velocity close to the speed of light to ensure sustained synchronization with the laser pulse. The recent concept of employing relativistic plasma mirrors as injectors offers an effective solution to both requirements Thévenet et al. (2016). When a p-polarized ultraintense laser obliquely irradiates a solid plasma mirror, surface electrons are periodically ejected at relativistic speeds via the relativistic oscillating mirror mechanism Bulanov et al. (1994); Lichters et al. (1996); Baeva et al. (2006); Thaury et al. (2007), naturally accumulating near the zero-crossing of the reflected laser’s electric field. Relativistic electron bunches with central energy reaching 10 MeV have been observed from plasma mirrors using 20-TW linearly polarized lasers Thévenet et al. (2016). Building upon this, subsequent studies have explored VLA using tightly focused radially polarized laser pulses, which generate a strong longitudinal electric field to drive acceleration Zaïm et al. (2017, 2020); Cao et al. (2021). To further enhance both the acceleration process and beam confinement, twisted lasers with a topological charge of , which exhibit both longitudinal electric and magnetic fields, have been proposed for generating monoenergetic attosecond electron bunches Shi et al. (2021).
Despite these advancements, a key challenge persists: electrons inevitably detach from the acceleration phase during the interaction. In conventional laser beams, the high-intensity region and peak intensity propagate at the speed of light in vacuum. However, since electrons cannot reach this speed, they inevitably fall behind the pulse envelope and exit the acceleration region. Additionally, the inherent divergence of conventional laser pulses beyond their focal waist, constrained by the Rayleigh range, leads to a decrease in peak intensity, thereby limiting efficient acceleration over long propagation distances.
In this letter, we propose VLA from plasma mirrors utilizing a newly developed technique, i.e., flying focus lasers Froula et al. (2019), to address the limitations of conventional lasers. By controlling the focal time or location of distinct portions of the laser pulse, the focal velocity of flying focus lasers can be decoupled from the group velocity of the pulse envelope, allowing it to propagate at any desired velocity Sainte-Marie et al. (2017); Froula et al. (2018); Jolly et al. (2020); Smartsev et al. (2019); Simpson et al. (2022); Ambat et al. (2023); Pigeon et al. (2023); Pierce et al. (2023). By setting the focal velocity to subluminal values, high-energy electrons can remain near the rising edge of the pulse envelope, within the longitudinal ponderomotive acceleration zone Ramsey et al. (2020, 2022), thereby staying in phase with the laser’s acceleration process. Moreover, flying focus beams can sustain near-constant peak intensities over longer distances compared to traditional Gaussian lasers, thus overcoming Rayleigh range limitations. These features can work synergistically to enhance the quality of electron beams in VLA. Using three-dimensional (3D) particle-in-cell (PIC) simulations, we demonstrate that flying focus lasers achieve a substantial increase, up to an order of magnitude, in the charge of relativistic electron bunches compared to conventional laser pulses.
II SIMULATION SETUP
The 3D PIC simulations are performed using the VLPL (Virtual Laser Plasma Lab) code Pukhov (1999). An -polarized flying focus beam is incident along the direction and focused onto an overdense plasma target at a 45∘ incidence angle. The reflected laser pulses propagate and accelerate the electrons injected from the plasma mirror along the direction (see Fig. 1(a)). The initial electric field of the flying focus pulse can be written as Ramsey et al. (2022, 2023):
where is the electric field amplitude at focus with the normalized laser vector potential , is the electron rest mass, is the speed of light in vacuum, is the elementary charge, is the beam waist, is the minimum beam waist, nm is the laser wavelength, with being the velocity of focus, with being the Rayleigh length, and are respectively the laser wave number and angular frequency, is the radial distance, with is a longitudinal super-Gaussian envelope to ensure the peak intensity of the flying focus beam maintain nearly constant during focal spot translation.
The fully ionized plasma has a density of and a thickness of 0.2, where is the critical density. The density decays exponentially with a scale length of at the front of the target. The ions are assumed to be immobile in the simulations. The simulation box has a size of . The grid step size in each direction is and the time step is . Absorption boundary condition is applied to both particles and fields. Limited by the significant computational time and storage requirements of 3D simulations, each cell is filled with 1 macroparticles. Nevertheless, the non-physical effects such as numerical heating are not significant. To track the electron dynamics far from the target, a moving window technique is employed, advancing at a speed of 0.9 along the + direction.
III RESULTS AND DISCUSSION
Figures 1(b) and (c) show the spatial distributions of the electric fields and the accelerated electrons with energy MeV, projected onto the plane, generated by conventional Gaussian and flying focus laser pulses at , respectively. Both lasers arrive at the target surface simultaneously, with identical peak intensities () and full width at half maximum (FWHM) pulse durations of . A train of femtosecond relativistic electron bunches is produced through VLA in both cases. However, a striking feature is that the number of electrons undergoing efficient acceleration and remaining synchronized with the laser field is significantly higher for flying focus lasers compared to Gaussian lasers. Additionally, the dephasing effect between the electrons and the laser field is evident, as some electrons are observed surfing in the deceleration phase. Figure 1(d) compares the energy spectra of all electrons at , when the laser peak first arrives at the plasma surface. It shows that the initial electron energy spectra generated by the two lasers are very similar. In fact, the high-energy tail of electrons produced by the Gaussian laser is even more pronounced than that from the flying focus laser. After the VLA process, however, the electron energy spectra at , shown in Fig. 1(e), indicate that the cutoff energy of the accelerated electrons driven by the flying focus lasers is now higher than that of the Gaussian lasers. More importantly, the number of relativistic electrons driven by the flying focus lasers is significantly greater–approximately one order of magnitude higher–than that produced by the Gaussian lasers. Specifically, the total charge of the accelerated electrons with energy MeV reaches 1.4 nC for the flying focus laser case, compared to only 0.2 nC for the Gaussian laser case.
To understand the differences in acceleration outcomes, we trace the evolution of electron dynamics during VLA. Figures 2(a) and (b) illustrate the initial positions of representative electrons shortly after their injection from plasma mirrors into the electric fields of Gaussian and flying focus lasers, respectively. Notably, the electron within the Gaussian laser field is in the acceleration phase, while the electron within the flying focus laser field is being decelerated. Figure 2(c) shows the electron’s - plane trajectory within the Gaussian laser field. The electron undergoes predominantly longitudinal acceleration for approximately 100, accompanied by transverse oscillations. However, the final energy gain remains low. Figures 2(d)–(f) display the temporal evolution of its transverse momentum , longitudinal momentum , and energy, respectively, all exhibiting oscillatory behavior. These oscillations, indicative of interaction with multiple Gaussian laser field optical cycles, arises from a lack of phase synchronization. The insufficient longitudinal momentum growth (Fig. 2(f)) prevents the electron from maintaining phase with the laser pulse, leading to repeated acceleration and deceleration, and ultimately restricting the net energy transfer. In contrast, the electron, despite initial deceleration in the flying focus laser field, shows a monotonic rise in momenta and energy after a single oscillation (see Figs. 2(g)–(j)). It is noteworthy that both multi-cycle interacting and phase-locked electrons are present in VLA for both Gaussian and flying focus beams. However, the crucial distinction is the significantly enhanced yield of phase-locked electrons in the flying focus beam’s VLA, driving its enhanced relativistic electron beam production. As depicted in Fig. 2(g), although the trajectory of the accelerated electron exhibits a continuous drift along the polarization direction, it remains spatially confined within the laser spot region. The observed increase in longitudinal momentum (Fig. 2(i)) is substantially greater than that of the transverse momentum (Fig. 2(h)), signifying that the electron experiences primarily longitudinal acceleration. Figure 2(k) shows the instantaneous work done by each field component as a function of time. It is evident that the longitudinal acceleration is predominantly driven by the combined contribution of the transverse magnetic and electric fields, highlighting the critical role of the laser’s longitudinal ponderomotive force.
We then examine the spatial profile evolution of the laser’s longitudinal ponderomotive force at various acceleration stages (Figure 3). Given the identical transverse Gaussian profiles of both flying focus and Gaussian lasers, as can be seen from Eq. (1), we analyze the one-dimensional longitudinal ponderomotive forces along the optical axis. Initially, the peak strength of the two lasers’ longitudinal ponderomotive force are comparable. However, the peak intensity of Gaussian lasers propagates at the speed of light , inevitably causing electrons to move out of the ponderomotive acceleration region and into the deceleration region. The phase-mismatched electrons interact with multiple laser periods, undergoing repeated cycles of acceleration and deceleration, which ultimately results in low energy gain. Additionally, due to diffraction over the Rayleigh length (approximately 78), the peak strength of the Gaussian lasers’ ponderomotive force decreases considerably. In contrast, the flying focus lasers maintain the peak strength of their longitudinal ponderomotive force over much longer distances, resulting in a higher cutoff energy for the accelerated electrons. Moreover, the peak intensity of the flying focus lasers propagates at a preset subluminal velocity of . As a result, a large population of high-energy electrons can remain within the ponderomotive acceleration region throughout the interaction, leading to a significantly larger number of accelerated electrons. These combined characteristics account for the superior acceleration performance observed with flying focus lasers.
Flying focus lasers can significantly enhance VLA electron charges across a broad range of laser-plasma parameters. Figure 4 shows the charge of accelerated electrons with energy MeV as a function of laser intensity. While Gaussian lasers also exhibit increased electron charge with intensity due to enhanced ponderomotive force and injection velocities, the flying focus laser demonstrates a steeper rise in electron charge. At an intensity level of approximately W/cm2, flying focus lasers achieve a total charge of approximately 10 nC for electrons with energies MeV. Such ultrashort relativistic electron bunches with substantial charge are highly advantageous for various applications, including single-shot electron radiography and the generation of high-flux x-rays Courtois et al. (2011); Bruhaug et al. (2023).
In addition to laser peak intensity, the focal velocity of the flying focus lasers also has an impact on the results. We have carried out simulations using different focal velocities and have also discovered significant enhancement in the charges of accelerated electrons. However, the expression for the flying focus beam employed in this study [Eq. (1)] inherently couples focal velocity adjustments to changes in pulse duration, which corresponds to the changes in the longitudinal ponderomotive force. This coupling makes it challenging to conduct an isolated study of the effects of focal velocity while maintaining a constant ponderomotive force.
IV Conclusion
In conclusion, 3D PIC simulations reveal that flying focus lasers dramatically improve the yield of high-energy electrons in the VLA scheme with plasma mirrors. This enhancement is driven by two primary characteristics. First, their subluminal intensity peak propagation allows for sustained synchronization between electrons and the laser field. Second, in comparison to a Gaussian beam, the flying focus beam maintains higher longitudinal ponderomotive force magnitudes over extended temporal and spatial scales. This approach offers superior relativistic electron sources for applications such as single-shot radiography and Thomson scattering, where high-flux beams are essential for achieving more accurate and detailed imaging and measurements.
acknowledgments
This work was supported by the National Natural Science Foundation of China (12175157).